# Escaping the Verifier: Learning to Reason via Demonstrations

### v1

@@ Abstract
Training Large Language Models (LLMs) to reason often relies on Reinforcement Learning (RL) with task-specific verifiers. However, many real-world reasoning-intensive tasks lack verifiers, despite offering abundant expert demonstrations that remain under-utilized for reasoning-focused training. We introduce RARO (Relativistic Adversarial Reasoning Optimization) that learns strong reasoning capabilities from only expert demonstrations via Inverse Reinforcement Learning. Our method sets up an adversarial interaction between a policy (generator) and a relativistic critic (discriminator): the policy learns to mimic expert answers, while the critic learns to compare and distinguish between policy and expert answers. Our method trains both the policy and the critic jointly and continuously via RL, and we identify the key stabilization techniques required for robust learning. Empirically, RARO significantly outperforms strong verifier-free baselines on all of our evaluation tasks — Countdown, DeepMath, and Poetry Writing — and enjoys the same robust scaling trends as RL on verifiable tasks. These results demonstrate that our method effectively elicits strong reasoning performance from expert demonstrations alone, enabling robust reasoning learning even when task-specific verifiers are unavailable.
@@ 1 Introduction
Recent advances in Large Language Models (LLMs) have been driven substantially by improvements in their reasoning abilities. Reasoning enables LLMs to perform deliberate intermediate computations before producing answers to the user queries, proposing candidate solutions and self-corrections. Much of this progress has been enabled via Reinforcemeng Learning (RL) on verifiable tasks such as mathematics and competitive programming (deepseek-r1; qwen3; grpo-deepseekmath; deepcoder2025). Notably, recent work has demonstrated that RL with Verifiable Rewards (RLVR) can enable LLMs to develop robust reasoning capabilities without any additional supervision  (deepseek-r1). A growing body of work further improves the efficiency and stability of such RL algorithms on verifiable tasks, such as DAPO (yu2025dapoopensourcellmreinforcement) and GSPO (gspo). However, comparatively little attention has been paid to developing reasoning abilities on non-verifiable tasks, where task-specific verifiers are unavailable.
@@
Without preference data, the typical approach to improving LLM performance in these domains is to conduct Supervised Fine-Tuning (SFT) on expert demonstration data via the next-token prediction objective. However, such methods, even if the data are further annotated with reasoning traces, does not encourage the same reasoning behaviors ellcited from large-scale RL training on verifiable tasks  (chu2025sftmemorizesrlgeneralizes). Additionally, naive next-token prediction objective induces training-inference distribution mismatch: during training, the model conditions only on the dataset contexts, whereas at inference, it conditions on self-sampled contexts. Training on self-sampled contexts, as occurs during RL, yields lower training-inference mismatch, leading to better performance at test time  (ross-learning-upper-bound). Thus, we hypothesize that leveraging expert demonstrations in conjunction with RL could cultivate robust reasoning abilities, leading to substantially improved performance on downstream tasks and offering a new pathway for developing reasoning capabilities in non-verifiable domains.
@@
  • We evaluate RARO on a controlled toy reasoning task, Countdown, where it not only significantly outperforms SOTA baselines without verification (zhou2025verifree), but it nearly matches the performance of RLVR, demonstrating the effectiveness of RARO on inducing reasoning behaviors.
@@
2 Preliminary
2.1 LLM Reasoning
Reasoning in LLMs has been a central focus of recent work, with numerous approaches proposed to enhance it, including Chain-of-Thought (CoT) prompting (wei2022chain), Tree of Thoughts (ToT) (yao2023tree), and Buffer of Thoughts (BoT) (yang2024bufferthoughtsthoughtaugmentedreasoning).
CoT prompting is a simple yet effective technique: it enables LLMs to generate intermediate reasoning tokens that steer them toward correct answers without additional training. CoT also pairs naturally with Test-Time Scaling, critical for further performance gains (snell2024scalingllmtesttimecompute). Recently, reasoning LLMs operationalize this idea by explicitly training via RL with verifiable rewards to produce long CoT reasonings before outputting the final response, yielding substantially higher-quality answers (deepseek-r1; qwen3).
While CoT reasonings provide useful guidance, the ultimate objective is to improve the quality of the final answers. Thus, following prior work (trice), we take the perspective that an LLM can be modeled as a joint distribution over prompts, CoT reasonings, and answers. This perspective induces a conditional latent-variable model where the prompt is the conditioning variable, the answer is observed, while the CoT reasonings are latent variables.
@@ 2.2 Reinforcement Learning with Verifiable Rewards
Reinforcement Learning with Verifiable Rewards (RLVR) is a family of methods designed to train LLMs to reason on verifiable tasks, such as mathematics and competitive programming, enabling recent SOTA open-source models to achieve expert-level performance on relevant benchmarks  (deepseek-r1; yang2025qwen3technicalreport).
The dominant method in this line of work is Group Relative Policy Optimization (GRPO)  (grpo-deepseekmath), which builds upon the Proximal Policy Optimization (PPO) (ppo) algorithm by replacing the advantage function with a sample average computed from rollout groups. Since the introduction of GRPO, numerous works have been proposed to further improve its training efficiency and stability, such as DAPO (yu2025dapoopensourcellmreinforcement) and GSPO (gspo).
@@ 2.3 General Reasoning Learning
Although no consensus method exists for general reasoning learning to our knowledge, several recent efforts make early progress. zhou2025verifree propose to train LLMs to reason with reward derived from the model’s own logits on expert answers rather than from an external verifier. jia2025writingzerobridgegapnonverifiable propose a pairwise generative reward model with a PPO-style objective for non-verifiable writing tasks, achieving sizable gains without external training signals. ma2025generalreasoneradvancingllmreasoning distill a model-based verifier from a strong teacher and leverage it as a reward model to train a general reasoner without verifiers. li2025internbootcamptechnicalreportboosting investigate large-scale multi-task RLVR, hypothesizing that breadth across many tasks induces stronger general reasoning. We build on this line of work while adopting a complementary perspective based on Inverse Reinforcement Learning.
@@ 2.4 Inverse Reinforcement Learning
Recently, IRL has been applied to LLMs. sun2025inverserlignmentlargelanguagemodel propose to align LLMs to human demonstrations via IRL without preference labels or explicit feedback, showing that IRL is effective for aligning LLMs to human preferences without preference data. However, their applications remain limited in scope, and with the rise of RL for reasoning, it remains open whether IRL-based methods can effectively train reasoning-focused LLMs.
@@ 3 Method
We study the general setting where we are given an expert Question–Answer (QA) dataset, and we aim to train a LLM policy to produce expert-level answers via explicit reasoning. We adopt this setting because verifiable tasks are relatively scarce, whereas expert demonstration data are abundant for many non-verifiable domains (e.g., highly upvoted Stack Exchange answers).
To approach this task, we propose a novel inverse reinforcement learning framework that sets up an adversarial interaction between a policy and a relativistic critic: the policy learns to mimic expert answers, while the critic learns to discriminate between policy and expert answers via pairwise comparison. By jointly training both the policy and the critic to reason via RL, we enable the emergence of strong reasoning capabilities from demonstrations alone, without requiring task-specific verifiers.
3.1 From Maximum Likelihood to Reward Gradient
@@
Let $D=\{(q_{i},a_{i})\}_{i=1}^{n}$ denote the expert QA dataset. We parameterize our LLM policy as $\pi_{\theta}(a,z\mid q)$, a distribution over answers $a$ and Chain-of-Thought (CoT) reasonings $z$ conditioned on a question $q$. We let $\hat{p}_{q}(\cdot)$ denote the empirical distribution of questions in $D$, $\hat{p}_{a\mid q}(\cdot\mid\cdot)$ denote the empirical distribution of expert answers conditioned on a question, and the joint $\hat{p}_{D}=\hat{p}_{a\mid q}\hat{p}_{q}$ denote the empirical distribution of dataset pairs $(q,a)$.
@@
To address this intractability, we adopt the perspective of Inverse Reinforcement Learning (IRL). Rather than maximizing the marginal likelihood directly, we learn a parameterized reward $r_{\phi}(a,q)$ over QA pairs such that optimizing a policy $\pi_{\theta}(a\mid q)$ with respect to $r_{\phi}$ yields a “near-optimal” policy that approximately maximizes the ML objective.
We formalize “near-optimality” via the KL-regularized reward-maximization objective, and under this objective, it can be shown (awr) that the optimal policy has the following closed-form solution:
@@
where $Z_{\theta^{\star}(\phi)}(q)$ is the partition function, $\pi_{\mathrm{ref}}$ is a fixed reference policy, and $\beta>0$ controls the strength of the KL-regularization. See Appendix A.1 for the proof.
Reward Gradient.
With the closed-form expression for the optimal policy under the reward model $r_{\phi}$, we can derive the corresponding gradient needed to optimize it by differentiating the negative ML loss with respect to $\phi$:
@@
See Appendix A.2 for the proof.
Intuitively, the gradient shapes the reward signal by increasing $r_{\phi}(a,q)$ on expert answers and decreasing it on policy answers, thus nudging the policy toward the expert distribution. A concrete algorithm for the alternating optimization is given in Algorithm 4 in Appendix E.
3.2 Reasoning Reward Model
While Algorithm 4 provides a concrete method for optimization, we have yet to decide on an appropriate architecture for the reward model $r_{\phi}(\cdot)$. Our setting targets difficult QA tasks that benefit from reasoning. Consequently, to reliably separate expert from policy answers, we expect that the reward model should be at least as capable as the policy. Thus, a natural instantiation is therefore a reasoning LLM. Specifically, we reparametrize the reward model with a binary classification setup:
$r_{\phi}(a,q)=c_{\phi}(\ell=\mathrm{expert}\mid a,q)\;-\;c_{\phi}(\ell=\mathrm{policy}\mid a,q)$
where $c_{\phi}(\cdot)$ is a reasoning critic that classifies whether an answer is from the expert or the policy.
Under this parameterization, as shown in Appendix A.3, the gradient $\nabla_{\phi}L$ corresponds to the standard policy gradient, and we can further derive an unbiased estimator for $r_{\phi}(a,q)$, resulting in two simple reward functions for the critic and policy.
@@ 3.2 Reasoning Reward Model
$R_{\text{critic}}(\ell,a,q)=\mathbbm{1}_{\ell\text{ is correct}}-\mathbbm{1}_{\ell\text{ is incorrect}}.$
@@
$R_{\text{policy}}(a,q)=\mathbbm{1}_{\ell=\mathrm{expert}}-\mathbbm{1}_{\ell=\mathrm{policy}},\quad\ell\sim c_{\phi}(\cdot\mid a,q).$
This allows us to optimize both the critic and policy using any reward-maximization algorithm (e.g., GRPO). Intuitively, such reward formulation creates an adversarial game between the critic and policy: the critic is rewarded when it correctly classifies answer as coming from the expert or policy, while the policy is rewarded when the critic incorrectly classifies its answer as an expert answer.
Limitations.
Despite the theoreical soundness, this binary classification setup poses challenges for critic learning. As policy approaches the expert distribution, the classification task becomes much more difficult due to a lack of reference answer for the critic to compare against. In addition, with a optimal policy, the critic effectively degenerates to random guessing, providing high-variance, uninformative gradients to the policy, leading to training instability as observed in our initial experiments (see Appendix D.2).
3.3 Relativistic Critic
To address the lack of reference in the binary classification setup, we adopt a relativistic formulation: the critic takes a triplet $(q,a,a^{*})$ consisting of one policy answer and one expert answer, and outputs which is better or tie if they are equal in quality. This resolves the degeneracy where the critic is forced to differentiate even when the policy is optimal. We empirically show that the tie option is crucial for better performance (see Appendix D.2).
@@ 3.3 Relativistic Critic
Intuitively, unlike the binary classification setup, the relativistic critic is now given a pairwise comparison task: the critic is rewarded when it correctly identifies the expert answer, and the policy is rewarded when the critic mistakenly identifies its answer as the expert answer, with additional tie rewards to ensure non-degeneracy and stable learning. Algorithm 1 describes the full training process and see Appendix E for example critic outputs.
Algorithm 1 Relativistic Critic with KL-Regularized Policy
Inputs: Dataset $D=\{(q_{i},a_{i})\}$; Tie reward $\tau_{\mathrm{pol}},\tau_{\mathrm{crit}}$; Batch $B$; Rollout $K_{\pi},K_{c}$;
Models: Policy $\pi_{\theta}(z,a\mid q)$; Relativistic critic $c_{\phi}(q,a^{(1)},a^{(2)})$.
1:Initialize $\theta,\phi$
2:for $t=1,\dots,T$ do
3:  Draw $\{(q_{i},a^{\mathrm{E}}_{i})\}_{i=1}^{B}\sim D$
4:  for $i=1..B,k=1..K_{\pi}$ do
[…]
6:  end for
7:  for $i=1..B,j=1..K_{c}$ do
[…]
9:  end for
10:  GRPO step on $\theta$ to maximize $\mathbb{E}[R^{\mathrm{pol}}_{i,k}]-\beta D_{\mathrm{KL}}(\pi_{\theta}\|\pi_{\mathrm{ref}})$.
11:  GRPO step on $\phi$ to maximize $\mathbb{E}[R^{\mathrm{crit}}_{i,j}]$.
12:end for
13:return $\theta,\phi$

### v3

@@ Abstract
Training Large Language Models (LLMs) to reason often relies on Reinforcement Learning (RL) with task-specific verifiers. However, many real-world reasoning-intensive tasks lack verifiers, despite offering abundant expert demonstrations that remain under-utilized for reasoning-focused training. We introduce RARO (Relativistic Adversarial Reasoning Optimization) that learns strong reasoning capabilities from only expert demonstrations via Inverse Reinforcement Learning. Our method sets up an adversarial game between a policy and a relativistic critic: the policy learns to mimic expert answers, while the critic aims to identify the experts among (expert, policy) answer pairs. Both the policy and the critic are trained jointly and continuously via RL, and we identify the key stabilization techniques required for robust learning. Empirically, RARO significantly outperforms strong verifier-free baselines on all of our evaluation tasks — Countdown, DeepMath, and Poetry Writing — and enjoys the same robust scaling trends as RL with verifiers. These results demonstrate that our method effectively elicits strong reasoning performance from expert demonstrations alone, enabling robust reasoning learning even when task-specific verifiers are unavailable.
@@ 1 Introduction
Recent advances in Large Language Models (LLMs) have been driven substantially by improvements in their reasoning abilities. Reasoning enables LLMs to perform deliberate intermediate computations before producing answers to the user queries, proposing candidate solutions and self-corrections. Much of this progress has been enabled via Reinforcement Learning (RL) on verifiable tasks such as mathematics and competitive programming (deepseek-r1; qwen3; grpo-deepseekmath; deepcoder2025). Notably, recent work has demonstrated that RL with Verifiable Rewards (RLVR) can enable LLMs to develop robust reasoning capabilities without any additional supervision  (deepseek-r1). A growing body of work further improves the efficiency and stability of such RL algorithms on verifiable tasks, such as DAPO (yu2025dapoopensourcellmreinforcement) and GSPO (gspo). However, comparatively little attention has been paid to developing reasoning abilities on non-verifiable tasks, where task-specific verifiers are unavailable.
@@
Without preference data, the typical approach to improving LLM performance in these domains is to conduct Supervised Fine-Tuning (SFT) on expert demonstration data via the next-token prediction objective. However, such methods, even if the data are further annotated with reasoning traces, does not encourage the same reasoning behaviors elicited from large-scale RL training on verifiable tasks  (chu2025sftmemorizesrlgeneralizes). Additionally, naive next-token prediction objective induces training-inference distribution mismatch: during training, the model conditions only on the dataset contexts, whereas at inference, it conditions on self-sampled contexts. Training on self-sampled contexts, as occurs during RL, yields lower training-inference mismatch, leading to better performance at test time  (ross-learning-upper-bound). Thus, we hypothesize that leveraging expert demonstrations in conjunction with RL could cultivate robust reasoning abilities, leading to improved performance on downstream tasks and offering a new pathway for developing reasoning capabilities in non-verifiable domains.
@@
  • We evaluate RARO on a controlled toy reasoning task, Countdown, where it not only significantly outperforms SOTA baselines without verification, but it nearly matches the performance of RLVR, demonstrating the effectiveness of RARO on inducing reasoning behaviors.
@@
Figure 1: Overview of RARO. The method creates an adversarial game between a policy and a relativistic critic that share the same weights. The critic is rewarded for identifying the experts among (expert, policy) answer pairs, while the policy is rewarded for deceiving the critic. Additionally, the critic can declare tie, yielding stable rewards when the it is unsure. Both the policy and the critic are trained jointly and continuously via RL.
2 Related Work
@@ 2.1 Reinforcement Learning with Verifiable Rewards
Chain-of-Thought (CoT) prompting (wei2022chain) is a simple yet effective technique that enables LLMs to generate intermediate reasoning tokens, steering them toward correct answers. This approach has become a central focus for enhancing reasoning, pairing naturally with test-time scaling for further performance gains (snell2024scalingllmtesttimecompute).
Currently, the standard method to elicit such strong CoT capabilities is Reinforcement Learning with Verifiable Rewards (RLVR). RLVR methods train LLMs to produce long reasoning traces on tasks with ground-truth verifiers, enabling recent open-source models to achieve expert-level performance (deepseek-r1; yang2025qwen3technicalreport).
The dominant algorithm in this space is Group Relative Policy Optimization (GRPO) (grpo-deepseekmath), which builds upon Proximal Policy Optimization (PPO) (ppo) by estimating advantages via group-wise sample averages. Subsequent works like DAPO (yu2025dapoopensourcellmreinforcement) and GSPO (gspo) have further improved its efficiency and stability. However, while RLVR is highly effective, it is fundamentally constrained to verifiable tasks such as mathematics and competitive programming.
@@ 2.2 General Reasoning Learning
Although no consensus method exists for general reasoning learning to our knowledge, several recent efforts make early progress. zhou2025verifree and gurung2025learningreasonlongformstory propose to train LLMs to reason with reward derived from the model’s own logits on expert answers rather than from an external verifier. jia2025writingzerobridgegapnonverifiable propose a pairwise generative reward model with a PPO-style objective for non-verifiable writing tasks, achieving gains without external training signals. gunjal2025rubricsrewardsreinforcementlearning propose using an LLM-as-judge (gu2025surveyllmasajudge) together with pre-generated rubrics from strong LLM to provide rewards for non-verifiable tasks. ma2025generalreasoneradvancingllmreasoning distill a model-based verifier from a strong teacher to train general reasoners without rule-based verifiers. li2025internbootcamptechnicalreportboosting investigate large-scale multi-task RLVR, hypothesizing that breadth across many tasks induces stronger general reasoning. We build on this line of work while adopting a demonstration-only setting and a complementary perspective based on Inverse Reinforcement Learning.
@@ 2.3 Inverse Reinforcement Learning
sun2025inverserlignmentlargelanguagemodel recently investigated the application of IRL for aligning LLMs with expert demonstrations. They show that a classifier trained in the IRL paradigm can serve as an effective reward model for Best-of-N sampling. However, their work stops short of exploring stable, joint adversarial training, or reasoning-intensive tasks, where the model must learn to navigate complex solution spaces rather than aligning with surface-level preferences.
@@ 3 Method
We study the general setting where we are given an expert Question–Answer (QA) dataset, and we aim to train a reasoning LLM policy to produce expert-level answers via explicit CoT reasoning. We adopt this setting because verifiable tasks are relatively scarce, whereas expert demonstration data are abundant for many non-verifiable domains (e.g., highly upvoted Stack Exchange answers).
To approach this task, we propose a novel inverse reinforcement learning framework that sets up an adversarial interaction between a reasoning policy and a relativistic critic: the policy learns to output expert-like answers, while the critic learns to discriminate between policy and expert answers via pairwise comparison. By jointly training both the policy and the critic to reason via RL, we enable the emergence of strong reasoning capabilities from demonstrations alone, without requiring task-specific verifiers.
3.1 Reasoning Optimization via Inverse Reinforcement Learning
@@
Let $D=\{(q_{i},a_{i})\}_{i=1}^{n}$ denote the expert QA dataset. We parameterize our reasoning policy as a conditional latent-variable model (trice) $\pi_{\theta}(a,z\mid q)$, a distribution over answers $a$ and CoT reasonings $z$ conditioned on a question $q$. We let $\hat{p}_{q}$ denote the empirical distribution of questions in $D$, $\hat{p}_{a\mid q}$ denote the empirical distribution of expert answers conditioned on a question, and the joint $\hat{p}_{D}=\hat{p}_{a\mid q}\hat{p}_{q}$ denote the empirical distribution of dataset pairs $(q,a)$.
@@
To address this intractability, we adopt the perspective of Inverse Reinforcement Learning (IRL). Rather than maximizing the marginal likelihood directly, we learn a parameterized reward $r_{\phi}(a,q)$ over QA pairs such that optimizing $\pi_{\theta}$ with respect to $r_{\phi}$ yields a “near-optimal” policy that approximately maximizes the ML objective.
We formalize “near-optimality” via the KL-regularized reward-maximization objective. Under this objective, it can be shown (awr) that the optimal policy has marginal:
@@
where $Z_{\theta^{\star}(\phi)}(q)$ is the partition function, $\pi_{\mathrm{ref}}$ is a fixed reference policy, and $\beta>0$ controls the strength of the KL-regularization. See Appendix A.1 for the full proof.
With a closed-form expression for the optimal marginal under reward model $r_{\phi}$, we have a concrete pathway to optimize $\phi$ such that $\pi_{\theta^{\star}(\phi)}$ maximizes the ML objective. Specifically, as shown in Appendix A.2, we can directly optimize $\phi$ via gradient ascent with the following gradient:
@@
Finally, approximating $\pi_{\theta^{\star}(\phi)}$ by optimizing $\pi_{\theta}$ with reward $r_{\phi}$ via RL, we have an algorithm that optimizes reasoning policy $\pi_{\theta}$ with respect to the ML objective as shown in Algorithm 1.
Algorithm 1 Alternating Policy-Reward Optimization
Models: Reward $r_{\phi}(a,q)$; Reasoning policy $\pi_{\theta}(a,z\mid q)$.
1:Initialize $\phi,\theta$
2:for $t=1,\ldots,T$ do
5:  Reward update:
6:  Policy update: KL-regularized GRPO with $r_{\phi}(a,q)$ as the reward.
7:end for
8:return $\theta,\phi$
3.2 Reasoning Critic as Reward Model
To instantiate this framework, we need to decide on an appropriate architecture for the reward model $r_{\phi}$. Our setting targets difficult QA tasks that benefit from reasoning. Thus, to reliably separate expert from policy answers, we expect the reward model to be at least as capable as the policy.
To this end, we represent $r_{\phi}$ with a reasoning critic $c_{\phi}$ that takes an $(q,a)$ pair and outputs $\ell\in\{\mathrm{expert},\mathrm{policy}\}$, classifying whether an answer is from the expert or the policy. Specifically, we parametrize $r_{\phi}$ as the probability of classifying as expert for a $(q,a)$ pair.
Under this parameterization, as shown in Appendix A.3, the gradient $\nabla_{\phi}L$ corresponds to the standard policy gradient, resulting in simple reward functions for the critic and policy.
@@ 3.2 Reasoning Critic as Reward Model
$R_{\text{critic}}(\ell,a,q)=\mathbbm{1}_{\ell\text{ is correct}}$
@@
$R_{\text{policy}}(a,q)=\mathbbm{1}_{\ell=\mathrm{expert}},\quad\ell\sim c_{\phi}(\cdot\mid a,q)$
This allows us to optimize both the critic and policy using GRPO.
Such reward formulation creates an adversarial game between the critic and policy: the critic is rewarded when it correctly classifies answer as coming from the expert or policy, while the policy is rewarded when the critic incorrectly classifies its answer as an expert answer.
3.3 Relativistic Critic: The Pairwise Advantage
Despite the theoretical soundness, the previous setup poses challenges for critic learning. As policy approaches the expert, the classification task becomes much more difficult due to a lack of reference answer for the critic to compare against. In addition, with an optimal policy, the critic effectively degenerates to random guessing, providing high-variance, uninformative gradients to the policy, leading to training instability as we observed (see Appendix D.2).
To address these limitations, we instead adopt a relativistic formulation: the critic takes a triplet $(q,a,a^{*})$ consisting of one policy answer and one expert answer, and outputs which is better or tie if they are equal in quality. This resolves the degeneracy where the critic is forced to differentiate even when the policy is optimal. We empirically show that the tie option is crucial for better performance (see Appendix D.2).
@@ 3.3 Relativistic Critic: The Pairwise Advantage
Unlike the binary classification setup, the relativistic critic is now given a pairwise comparison task: the critic is rewarded when it correctly identifies the expert answer, and the policy is rewarded when the critic mistakenly identifies its answer as the expert answer, with additional tie rewards to ensure non-degeneracy and stable learning.